OpenStudy (anonymous):

What are the sine, cosine, and tangent of 5 pi over 4 radians?

OpenStudy (anonymous):

@hartnn @zaibali.qasmi

OpenStudy (anonymous):

see 1st convert it into degree form..it will be more easy to solve for you.. so 5*180/4=900/4=225 now 225=180+45 so use sin(180+45)=sin180cos45+cos180sin45 cos(180+45)=cos180cos45+sin180sin45 and tan(180+45)=\[\frac{ \tan180\tan45 }{ 1+\tan180\tan45 }\]

OpenStudy (anonymous):

simply plug in values from table and get result///

OpenStudy (anonymous):

? sorry im confused

OpenStudy (anonymous):

so what?

OpenStudy (anonymous):

@zaibali.qasmi i dont know how to do this...

OpenStudy (anonymous):

sin180=0 sin45=1/sqrt2 cos180=-1 cos45=1/sqrt2 tan180=0 tan45=1 put these values and solve..

OpenStudy (anonymous):

put them into what?